You are planning to invest $1,000 in an account earning 8​% per year for retirement.

a. If you put the $1,000 in an account at age​ 23, and withdraw it 44 years​ later, how much will you​ have?

b. If you wait 10 years before making the​ deposit, so that it stays in the account for only 34 years, how much will you have at the​ end?
Answer #1
a) Amount in the account after 44 years = 1000*1.08^ 44 = $ 29,555.97
b) Amount in the account after 34 years = 1000*1.08^ 34 = $ 13,690.13